# Learned Indexes with Distribution Smoothing via Virtual Points

### v1

@@ IV CDF Smoothing for a Single Linear Model
We start with a single indexing function over a segment of keys $K_{i}$. The optimisation goal can be written as:
$\text{argmin}_{V_{i},\theta}\mathcal{L}_{f_{\theta}}(K_{i}\cup V_{i}),\quad s.t.|V_{i}|\leq\lambda$ (3)
Here, $V_{i}$ represents the set of virtual keys added to the key set, and $\theta$ denotes the parameters used to define the refitted indexing function $f$ over $K_{i}\cup V_{i}$ – recall that we refit the indexing function after virtual point insertions. We include $V_{i}$ in the calculation to fit also the virtual points, such that the virtual points can be used to accommodate future insertions.
Solution overview. To locate the best virtual point to be inserted, we need to know the new SSE $\mathcal{L}_{f_{\theta}}(K_{i}\cup V_{i})$ after inserting each potential virtual point. To calculate the SSE for just one virtual point, it takes $O(n)$ time where $n$ is the size of $K_{i}$. Suppose that there are $p$ possible values (and hence $p$ candidate insertion positions) for the virtual point. Then, $O(n\cdot p)$ time is needed to find the best value for the insertion.
As both $n$ and $p$ can be large for real datasets, we propose an efficient solution with three main steps: (1) We calculate the loss function value of each candidate insertion position (that is, what will be the error if a point is inserted at that position; Section IV-A). (2) Based on the calculated loss function values and their first partial derivatives, we filter the candidate positions (Section IV-B). (3) From the remaining candidates, we present an efficient algorithm to find the best subset of virtual points of size $\lambda$ (Section IV-C).
@@
Figure 3: Prediction errors (SSE) corresponding to different insertion positions for a virtual point.
To reduce the search space for the virtual point insertions, we bound it between $(\min\{K_{i}\},\max\{K_{i}\})$. This is because any virtual points added prior to $\min\{K_{i}\}$ would cause all keys’ ranks to increase at the same time, while adding virtual keys after $\max\{K_{i}\}$ would not impact any key’s rank. As such, neither would help achieve a better-fitted indexing function. We also skip the index keys already in $K_{i}$, such that our solution can be compatible with learned indexes that do not support duplicate keys [21]. These skipped keys break the candidate virtual point values (i.e., insertion positions) into segments (i.e., sub-sequences) as shown in Fig. 3, which corresponds to the keys in Fig. 2(a). Each of such sub-sequences are convex, as shown in [29]. In Fig. 3, every hollow dot represents a candidate virtual point. Its $y$-value represents the new loss value (i.e., SSE) when the virtual point is included into the key set. Adjacent hollow dots are linked together, forming a segment of virtual point values. For example, the segment formed by 21 to 25 in the figure is between index keys 20 and 26. Inserting a virtual point with value 23 leads to the smallest SSE as shown in the figure.
IV-A Inserting One Virtual Point
We first optimise the loss calculation for each candidate virtual point. It is computationally expensive to calculate the SSE for each candidate virtual point due to the size of the original key set. We present the following modified equations for the refitting of the index function, such that the included virtual point is separated from the values in the original key set, making the equations dependent only on the inserted virtual point. This way, we enable a single total SSE calculation for the original key, which is then reused to reduce the computational costs when computing the total SSE for the candidate virtual points.
Equations 4 and 5 below compute the new coefficients ($w$ and $b$) for the indexing function $f_{i}$ (shown in Fig. 2(b)).
$w=\frac{(\sum_{i=1}^{n}k_{i}y_{i}+k_{v}y_{v})-(n+1)\bar{k_{v}}\bar{y_{v}}}{(\sum_{i=1}^{n}k_{i}^{2}+k_{v}^{2})-(n+1)\bar{k_{v}^{2}}},$ (4)
$b=\bar{y_{v}}-w\bar{k_{v}}.$ (5)
@@ IV-A Inserting One Virtual Point
Here, the new virtual point is given by $k_{v}$, and the rank of key $k_{i}$ is given by $y_{i}$ (assuming that the ranks start from 0). Further, $\bar{k_{v}}$ and $\bar{y_{v}}$ are the mean of the key set and the rank set after inserting the virtual point $(k_{v},y_{v})$, respectively (i.e., the means of the keys and ranks in Fig. 2(b)). They can be computed by Equations 6 and 7 as follows:
$\bar{k_{v}}=\frac{\sum_{i=1}^{n}k_{i}+k_{v}}{n+1},$ (6)
$\bar{y_{v}}=\frac{\sum_{i=1}^{n}y_{i}+n}{n+1}.$ (7)
Equations 8 to 11 are devised for the efficient computation of the terms in Equations 4 and 5.
$\sum_{i=1}^{n}y_{i}=\sum_{i=1}^{n}y_{original_{i}}+n-y_{v},$ (8)
$\bar{y}=\frac{\sum_{i=1}^{n}y_{i}}{n},$ (9)
$\bar{k}=\frac{\sum_{i=1}^{n}k_{i}}{n},$ (10)
$\sum_{i=1}^{n}k_{i}y_{i}=\sum_{i=1}^{n}k_{i}y_{original_{i}}+\sum_{i=y_{v}}^{n}k_{i}.$ (11)
@@
Here, $\bar{k}$ and $\bar{y}$ are the mean of the key set and the rank set before inserting the virtual point $(k_{v},y_{v})$, respectively (i.e., the means of the keys and ranks in Fig. 2(a)); $y_{original_{i}}$ refers to the rank of key $k_{i}$ prior to inserting the virtual point.
The prediction error function, where the virtual point is separated, is given in Equation 12. This rewritten function enables the values for any candidate virtual point to be calculated efficiently by simply adding the error of the candidate virtual point to the already calculated errors of the original key set. In the case of inserting $\lambda$ virtual points, after inserting one virtual point $(k_{v1},y_{v1})$ and then to find the next virtual point $(k_{v2},y_{v2})$, errors of the original key set will be adjusted to include the error of $(k_{v1},y_{v1})$. As a result, the total SSE values for the candidate virtual points for $(k_{v2},y_{v2})$ can also be easily calculated.
$\begin{split}\mathcal{L}_{i}(\{K_{i}\cup V_{i}\})=w^{2}\sum_{i=1}^{n}k_{i}^{2}+2wbn\bar{k}\\ -2w\sum_{i=1}^{n}k_{i}y_{i}+nb^{2}-2nb\bar{y}\\ +\sum_{i=1}^{n}y_{original_{i}}^{2}+n^{2}-y_{v}^{2}+(w{k_{v}}+b-y_{v})^{2}\end{split}$ (12)
IV-B Filtering Virtual Point Candidates
@@
In this subsection, we present an efficient approach to find the candidate virtual points that will reduce the loss function (i.e., total SSE) using the derivative of the loss function, thus providing a much smaller set of candidate virtual points. This step is important because while using the equations above helps improve the efficiency of processing one candidate virtual point, the number of candidate virtual points to be processed can still be large, causing substantial overheads.
@@
Fig. 4 plots the partial derivative of the total SSE with respect to a candidate virtual point. The sub-sequences (depicted as lines or dots) that cross the zero $x$-axis would imply that there is a minimum point within the sub-sequence. Otherwise, the minimum point would be in one of the endpoints of the sub-sequence of candidate virtual points. This observation is exploited to streamline the selection of candidate virtual points, i.e., to select the best virtual point from each sub-sequence.
Similar to the computation of the total SSE, the partial derivative of total SSE with respect to candidate virtual point $k_{v}$ can also be computed by Equation 13:
$\begin{split}{\mathcal{L}_{i}(\{K_{i}\cup V_{i}\})}^{\prime}=2({w}^{\prime}(w\sum_{i=1}^{n}k_{i}^{2}+nb\bar{k}-\sum_{i=1}^{n}k_{i}y_{i})+\\ n{b}^{\prime}(w\bar{k}+b-\bar{y})+(wk_{v}+b-y_{v})({w}^{\prime}k_{v}+w+{b}^{\prime})).\end{split}$ (13)
@@
Here, $w^{\prime}$ and $b^{\prime}$ refer to the partial derivative of the slope $w$ and the intercept $b$ with respect to the candidate virtual point $k_{v}$. They can be computed by Equations 14 and 15, respectively.
${w}^{\prime}=\frac{A(n(y_{v}-\bar{y}))-B(2n(k_{v}-\bar{k}))}{A^{2}},$ (14)
${b}^{\prime}=-\frac{(w+(n+1)\bar{k_{v}}{w}^{\prime})}{n+1},$ (15)
$A=(n+1)(\sum_{i=1}^{n}k_{i}^{2}+k_{v}^{2})-((n+1)\bar{k_{v}})^{2},$ (16)
$B=(n+1)(\sum_{i=1}^{n}k_{i}y_{i}+k_{v}y_{v})-(n+1)^{2}\bar{k_{v}}\bar{y_{v}}.$ (17)
@@
Figure 4: First partial derivative of the loss function with respect to different insertion positions for a virtual point
IV-C Inserting $\lambda$ Virtual Points
@@
When there are $\lambda$ virtual points to insert, the optimal solution would require computing the total SSE for every size-$\lambda$ subset of the candidate virtual points in the range of $[\min\{K_{i}\},\max\{K_{i}\}]$. If there are $p$ possible insertion positions for the virtual points, the time complexity will be $O(^{p}C_{\lambda}\cdot n\cdot p)$, where ${}^{p}C_{\lambda}$ is the combination of every size-$\lambda$ subset from $p$. As this will be prohibitively expensive for a large dataset, we propose a greedy algorithms and insert individual virtual points iteratively.
The core idea is to identify the virtual point that would minimise the total SSE for each sub-sequence (i.e., local minima for the sub-sequences) and select the one that reduces the total SSE the most (i.e., global minimum). This process needs to be performed $\lambda$ times. The algorithm for CDF smoothing by inserting $\lambda$ virtual points is summarized in Algorithm 1 and described below.
@@
1:Key set: $K$, loss function with new virtual point: $\mathcal{L}(k_{v})$, smoothing threshold: $\alpha$
@@
4:$\mathcal{L}^{\prime}(k_{v})=\frac{\partial L(k_{v})}{\partial k_{v}}$, $\lambda=\alpha\cdot K.size$, $\mathcal{L}_{previous}=\mathcal{L}(\varnothing)$
@@
15:         if $\mathcal{L}^{\prime}(G[i].first)\cdot\mathcal{L}^{\prime}(G[i].second)<$then
@@
The algorithm takes as input a key set $K$ (we drop the subscript $i$ from $K_{i}$ to simplify the notation) of size $n$ and a smoothing threshold $\alpha$ (or a smoothing budget $\lambda$ can also be directly provided ). We use $G$ to denote a set that stores the potential sub-sequences, where there can be a candidate virtual point within the sub-sequences with a local minima of the total SSE. The candidate virtual points contributing the local minima are stored in an array $M$, while $C$ stores set of candidate keys for the virtual points. We use $U$ to hold the total SSE value for each candidate virtual point and Vector $V$ to store the final optimal virtual points.
First, the algorithm identifies the sub-sequences of candidate virtual points that could have their minimal total SSE values at the endpoints or in the middle of the sequence. This is shown in Lines 4 to 12. If there are more than two points in a sub-sequence, the candidate virtual point with the minimal total SSE value can be within that sub-sequence. As such, the two endpoints of the sub-sequence are saved in array $G$ for calculating the partial derivatives.
Afterwards, in Lines 13 to 22, the partial derivative of the total SSE with respect to the candidate virtual points is calculated for all point pairs in $G$ using the equations derived above. If the signs of the partial derivatives corresponding to the two endpoints of a sub-sequence are different (i.e., on opposite sides of the $x$-axis), the two endpoints will be added to array $M$ for calculating the minimum point. As shown in Fig. 4, for the sub-sequences that contain candidate virtual points with minimal total SSE values, the partial derivatives of the end points will appear on the two sides of the $x$-axis. These minimal points are added to array $C$ after they are calculated. If the partial derivatives of the two endpoints are on the same side of the $x$-axis, the minima is at one of the endpoints, as such the two end points are added to $C$.
Finally, Lines 23 to 31 compute the total SSE for each candidate virtual point in $C$ and select the point with the minimum total SSE, as long as the new total SSE is smaller than the existing total SSE obtained so far over $K$ and any previously inserted virtual points. This process is repeated until at most $\lambda$ virtual points are inserted, or when the total SSE is not reduced any further. When the algorithm terminates, the final virtual points in $V$ are returned.
Complexity analysis. Our proposed CDF smoothing algorithm reduces the computation of the total SSE values over $K$ to just once, which takes $O(n)$ time. This process is repeated to find $\lambda$ optimal candidate virtual points. However, there is no need to recalculate the loss function value after adding a virtual point, as we could treat the key set with the previous virtual point inserted as the new original or base key set for a constant time calculation. Thereby, giving a time complexity of $O(\lambda+n)$.
V CDF Smoothing for Hierarchical Indexes

### v2

@@ 4. CDF Smoothing for a Single Linear Model
We first consider a single indexing function over a segment of keys $K_{i}$. Given the smoothing budget $\lambda$, our optimisation goal is:
(4) $\text{argmin}_{V_{i},w,b}\mathcal{L}_{f_{w,b}}(K_{i}\cup V_{i}),\quad s.t.|V_{i}|\leq\lambda$
This optimisation goal varies from Equation 2. Importantly, we now allow the slope ($w$) and the intercept ($b$) parameters of the indexing function $f$ to be refitted based on the keys $K_{i}$ (with adjusted ranks) and the inserted virtual points $V_{i}$, rather than just inserting virtual points to adjust $rank(k)$ for $k\in K_{i}$ to fit $f(k)$ of the original indexing function $f$. This way, intuitively, we could achieve a lower loss with fewer virtual point insertions (hence reducing space overhead), as opposed to the naive optimal smoothing scheme described above that simply spreads $K_{i}$ to fit $rank(k)$ to a given $f(k)$.
In Equation 4, we include $V_{i}$ in the loss calculation, such that the storage space allocated to the virtual points can be used to accumulate data insertions, with minimized prediction errors when querying the inserted data points.
Challenges. Allowing to refit the slope and the intercept of the indexing function makes the optimisation problem even more difficult, as now the loss reduction brought by inserting a virtual point depends further on the other virtual points to be inserted. To solve the problem, a simple greedy heuristic is to iteratively select the best remaining candidate virtual point that leads to the largest loss reduction and to refit the indexing function after each virtual point selection. This process is repeated for $\lambda$ iterations to identify all $\lambda$ virtual points to be inserted.
There are complexity issues with this simple heuristic.
Challenge 1: Large size of the candidate set. Given keys $K_{i}$, the candidate virtual points can be any key value in $(\min\{K_{i}\},\max\{K_{i}\})$ (detailed in Section 4.2), which can be a large range in real datasets.
Challenge 2: Repeated loss calculations. For each candidate virtual point $k_{v}$, we need to recalculate $\mathcal{L}_{f}(K_{i}\cup V_{i})$, where $V_{i}$ includes $k_{v}$, to help select the best candidate virtual point that leads to the largest loss reduction, which takes $O(|K_{i}|+\lambda)$ time. For $\lambda$ iterations and $O(p)$ candidate virtual points to be examined per iteration, overall, the greedy solution above takes $O((|K_{i}|+\lambda)\cdot\lambda\cdot p)$ time.
To overcome these challenges, we present an efficient solution below to reduce the overall time complexity through (1) reducing the number of candidate virtual points $p$ and (2) the time cost to calculate the loss for each candidate virtual point.
Our solution. Our solution takes three steps:
Step 1: We propose an effective approach to reduce the number of candidate virtual points. The idea is that, if there are consecutive candidate virtual points, we only need to consider the point among these consecutive points that minimises the loss function, which can be identified utilising the first-order partial derivative of the loss function (addressing Challenge 1, detailed in Section 4.2).
Step 2: We propose an efficient algorithm to calculate the loss. The core idea is to calculate the loss incrementally, and to reuse part of the calculation from the previous iteration, as only one new candidate virtual point is included into the calculation each time (addressing Challenge 2, detailed in Section 4.1).
Step 3: Finally, from the reduced set of candidate virtual points, and with efficient loss calculation, we present an efficient algorithm to compute the best subset of virtual points of size $\lambda$ (Section 4.3).
A running example on finding the best virtual point is shown in Fig. 3, which corresponds to the keys in Fig. 2(a). In Fig. 3, each hollow dot represents a candidate virtual point. Its $y$-value represents the loss (i.e., SSE) if the virtual point is included into the key set. Adjacent hollow dots are linked together, forming a segment of candidate virtual point values. For example, the segment formed by 21 to 25 is between index keys 20 and 26 (the index keys themselves are not considered to be candidate virtual points; they correspond to the gaps between the curve segments in Fig. 3). The goal is to search for the virtual point with the smallest loss value. In the given example, point 23 is the search target.
In what follows, we present the algorithm to efficiently calculate the loss first (Section 4.1), based on which the strategy to reduce the candidate virtual points is described (Section 4.2), and our overall algorithm to compute the best $\lambda$ virtual points is detailed (Section 4.3).
@@
Figure 3. Loss function (SSE) value corresponding to different insertion positions for a virtual point.
4.1. Efficient Loss Calculation and Indexing Function Refitting
We start with efficiently calculating the loss for each candidate virtual point, where the idea is to reuse the calculations as much as possible. For the rest of Section 4, we abuse the notation slightly and use $K$ instead of $K_{i}$ to denote the segment of keys for which CDF smoothing is to be done, since the discussion only concerns a single segment. Similarly, we use $f$ instead of $f_{i}$ to denote the indexing function, $\mathcal{L}$ instead of $\mathcal{L}_{i}$ to denote the loss function, and $V$ instead $V_{i}$ to denote the candidate virtual points for the segment.
Let the rank of a key $k_{i}$ be $y_{i}$ (assuming that the ranks start from 0) for the $n$ keys in $K$, i.e., $y_{i}=rank(k_{i})$. First, consider a virtual point $k_{v}$ with its rank (after inserted into $K$) being $y_{v}$ (we explain the case for multiple virtual points later). The loss for $K$ and $V=\{k_{v}\}$ can be calculated as follows, where $w$ and $b$ are the parameters of the indexing function $f$:
(5) ($\mathcal{L}(\{K\cup V\})=\sum_{i=1}^{n}(wk_{i}+b-y_{i})^{2}+(wk_{v}+b-y_{v})^{2}$
It is computationally expensive to calculate the loss for each candidate virtual point using Equation 5, as $w$ and $b$ can change when the indexing function $f$ is refitted for the candidate virtual point, which is governed by the following equations (derived from the first-order partial derivatives of Equation 5):
(6) $w=\frac{\sum_{i=1}^{n+1}(k_{i}-\bar{k_{v}})(y_{i}-\bar{y_{v}})}{\sum_{i=1}^{n+1}(k_{i}-\bar{k_{v}})^{2}},$
(7) $b=\bar{y_{v}}-w\bar{k_{v}}.$
@@ 4.1. Efficient Loss Calculation and Indexing Function Refitting
Here, $\bar{k_{v}}$ and $\bar{y_{v}}$ are the mean of the key set (i.e., $K\cup V$) and the rank set after inserting the virtual point $(k_{v},y_{v})$, respectively. They can be computed by Equations 8 and 9 as follows:
(8) $\bar{k_{v}}=\frac{\sum_{i=1}^{n}k_{i}+k_{v}}{n+1},$
(9) $\bar{y_{v}}=\frac{\sum_{i=1}^{n}y_{i}+n}{n+1}.$
Next, we rewrite Equation 5 such that the candidate virtual point $k_{v}$ is separated from the values of $K$. This enables us to calculate the terms in the equation related to $K$ separately and then reuse their values for different candidate virtual points. This would reduce the time when computing the loss for different candidate virtual points.
(10) ($\begin{split}\mathcal{L}(\{K\cup V\})=w^{2}\sum_{i=1}^{n}k_{i}^{2}+2wbn\bar{k}\\ -2w\sum_{i=1}^{n}k_{i}y_{i}+nb^{2}-2nb\bar{y}\\ +\sum_{i=1}^{n}y_{original_{i}}^{2}+n^{2}-y_{v}^{2}+(w{k_{v}}+b-y_{v})^{2},\end{split}$
where
(11) $\sum_{i=1}^{n}y_{i}=\sum_{i=1}^{n}y_{original_{i}}+n-y_{v},$
(12) $\bar{y}=\frac{\sum_{i=1}^{n}y_{i}}{n},$
(13) $\bar{k}=\frac{\sum_{i=1}^{n}k_{i}}{n},$
(14) $\sum_{i=1}^{n}k_{i}y_{i}=\sum_{i=1}^{n}k_{i}y_{original_{i}}+\sum_{i=y_{v}}^{n}k_{i}.$
@@
In the equations, $\bar{k}$ and $\bar{y}$ are the mean of the key set $K$ and the corresponding rank set before inserting the virtual point $(k_{v},y_{v})$, respectively; $y_{original_{i}}$ refers to the rank of key $k_{i}$ prior to inserting the virtual point.
Given the rewritten loss, Equations 15 and 16 are derived by separating the terms related to the candidate virtual point from those related to the original key set $K$, which enable more efficient computation of $w$ and $b$.
(15) $w=\frac{(\sum_{i=1}^{n}k_{i}y_{i}+k_{v}y_{v})-(n+1)\bar{k_{v}}\bar{y_{v}}}{(\sum_{i=1}^{n}k_{i}^{2}+k_{v}^{2})-(n+1)\bar{k_{v}^{2}}},$
(16) $b=\bar{y_{v}}-w\bar{k_{v}}.$
Adjustment for multiple virtual points. In the case of inserting $\lambda$ virtual points, after inserting one virtual point $(k_{v1},y_{v1})$, to find the next virtual point $(k_{v2},y_{v2})$, the original key set terms will be adjusted to include the newly added virtual point $(k_{v1},y_{v1})$. As a result, the loss for the next candidate virtual point $k_{v2}$ can also be efficiently calculated by considering only the changes induced by $k_{v2}$ and its corresponding $y_{v2}$.
In the derived equations above, the terms are separated based on whether they belong to the original key set or dependent on the candidate virtual key to be inserted. Doing so enables efficient calculations of the loss by reusing the terms of the original key set after calculating them just once.
4.2. Filtering Virtual Point Candidates
@@
Next, we present an efficient approach to identify the candidate virtual points that can potentially reduce the loss, thus providing a much smaller set of candidate virtual points to consider. This step is important because while using the equations above helps improve the efficiency of processing one candidate virtual point, the number of candidate virtual points to be processed has a multiplicative impact to the overall algorithm time efficiency.
To reduce the search space for the candidate virtual points, we bound it in $(\min\{K\},\max\{K\})$. This is because any virtual points added prior to $\min\{K\}$ would cause all keys’ ranks to increase at the same time, while adding virtual points after $\max\{K\}$ would not impact any key’s rank. As such, neither would help achieve a better-fitted indexing function. We also skip the index keys already in $K$, such that our solution can be compatible with learned indexes that do not support duplicate keys (Sun et al., 2023).
@@
Below, we present an approach based on the derivative of the loss function to further reduce the set of candidate virtual points. Our idea is illustrated using Fig. 4, which plots the partial derivative of the loss function with respect to a candidate virtual point $k_{v}$. Each sub-sequence (depicted as lines or dots) corresponds to the partial derivative of the loss of a sub-sequence of key values of a candidate virtual point (which can also be seen as a sub-sequence of candidate virtual points). The sub-sequences that cross the zero-value line (i.e., the $x$-axis) imply that there is a minimal loss point within the sub-sequence. Otherwise, the minimal point for the sub-sequence must be at one of two endpoints of the sub-sequence, since each of such sub-sequences has been shown to be convex (Kornaropoulos et al., 2022). Intuitively, this convex property is because the loss function is a summation of $n$ quadratic terms ($(wk_{i}+b-y_{i})^{2}$) and only one non-quadratic term ($(wk_{v}+b-y_{v})^{2}$), and the $wk_{v}$ term is nonlinear variable. We exploit this property to streamline the selection of candidate virtual points, i.e., to select the best virtual point from each sub-sequence.
Following the idea above, we propose to filter the candidate virtual points as follows:
  1. (1)
    For each sub-sequence of candidate points (that is, where the candidate virtual point values are continuous), if the length of the sub-sequence is greater than 2, there can be a candidate virtual point within the sub-sequences with a local minima of the loss. We compute the partial derivative of the loss function, which includes the previously added virtual points (($\mathcal{L}(\{K\cup V\})$ shown in Equation 5) with respect to candidate virtual point $k_{v}$, i.e., (${\mathcal{L}(\{K\cup V\})}^{\prime}$ of the two endpoints of such sub-sequence (we present efficient ways to compute the partial derivative of a candidate later).
    • If the sign of (${\mathcal{L}(\{K\cup V\})}^{\prime}$ of the two endpoints are the same, it means that there is no point with the local minima within this sub-sequence (i.e.,, the sub-sequence does not cross the zero value in $y$-axis as shown in Fig. 4). In that case, we only need to consider the two endpoints of the sub-sequence as the candidate virtual points, and we can safely discard all the candidate virtual points in between.
    • Otherwise, if the sign of (${\mathcal{L}(\{K\cup V\})}^{\prime}$ of the two endpoints are opposite, it means that there is a point with a local minima within the sub-sequence. The minimal point can be calculated by using the two partial derivative values to find their intersection with the x-axis. As this minimal point is guaranteed to have a smaller loss than all the other points in that sub-sequence, only the point is considered as a candidate virtual point. All other points in the sub-sequence can be safely discarded.
  2. (2)
    If the length of the sub-sequence is less than or equal to 2, we need to consider all points in the sub-sequence as the candidate virtual points.
Efficient computation of the first-order derivative of the loss function value. Similar to the computation of the loss function, we present an efficient way to calculate the first-order partial derivative of the loss function with respect to candidate virtual point $k_{v}$, where $k_{v}$ and its related terms are separated from the other terms to enable reusing the terms that require information from $K$ only. This is achieved by the following equation:
(17) ($\begin{split}{\mathcal{L}(\{K\cup V\})}^{\prime}=2({w}^{\prime}(w\sum_{i=1}^{n}k_{i}^{2}+nb\bar{k}-\sum_{i=1}^{n}k_{i}y_{i})+\\ n{b}^{\prime}(w\bar{k}+b-\bar{y})+(wk_{v}+b-y_{v})({w}^{\prime}k_{v}+w+{b}^{\prime})).\end{split}$
@@
Here, $w^{\prime}$ and $b^{\prime}$ refer to the partial derivatives of $w$ and $b$ with respect to $k_{v}$, respectively. They can be computed by Equations 18 and 19 as follows.
(18) ${w}^{\prime}=\frac{A(n(y_{v}-\bar{y}))-B(2n(k_{v}-\bar{k}))}{A^{2}},$
(19) ${b}^{\prime}=-\frac{(w+(n+1)\bar{k_{v}}{w}^{\prime})}{n+1},$
(20) $A=(n+1)(\sum_{i=1}^{n}k_{i}^{2}+k_{v}^{2})-((n+1)\bar{k_{v}})^{2},$
(21) $B=(n+1)(\sum_{i=1}^{n}k_{i}y_{i}+k_{v}y_{v})-(n+1)^{2}\bar{k_{v}}\bar{y_{v}}.$
@@
Figure 4. First-order partial derivatives of the loss (Equation 17) with respect to the key value of a virtual point $k_{v}$
4.3. Algorithm for Inserting $\lambda$ Virtual Points
@@
When there are $\lambda$ virtual points to insert, the optimal solution would require computing the loss for every size-$\lambda$ subset of the candidate virtual points in the range of $(\min\{K\},\max\{K\})$. If there are $p$ possible insertion positions for the virtual points, the time complexity will be $O(^{p}C_{\lambda}\cdot n\cdot p)$, where ${}^{p}C_{\lambda}$ is the combination of every size-$\lambda$ subset from $p$. As this will be prohibitively expensive for a large dataset, we propose a greedy algorithm that inserts individual virtual points iteratively.
Our core idea is to identify the virtual point that would minimise the loss for each sub-sequence (i.e., local minima for the sub-sequences) and select the one that reduces the loss the most (i.e., global minimum). This process needs to be performed $\lambda$ times. The algorithm for CDF smoothing by inserting $\lambda$ virtual points is summarised in Algorithm 1 and described below.
@@
1:Key set: $K$, loss function with new virtual point: $\mathcal{L}(K\cup k_{v})$, smoothing threshold: $\alpha$
@@
4:$\mathcal{L}^{\prime}(K\cup k_{v})=\frac{\partial L(K\cup k_{v})}{\partial k_{v}}$, $\lambda=\alpha\cdot K.size$, $\mathcal{L}_{previous}=\mathcal{L}(K\cup\varnothing)$
@@
15:         if $\mathcal{L}^{\prime}(K\cup G[i].first)\cdot\mathcal{L}^{\prime}(K\cup G[i].second)<$then
@@
The algorithm. Our algorithm takes as input a key set $K$ and a smoothing threshold $\alpha$ (or a smoothing budget $\lambda=\alpha n$). We use $G$ to denote a set that stores the potential sub-sequences, where there can be a candidate virtual point within the sub-sequences with a local minima of the loss. The candidate virtual points contributing the local minima are stored in an array $M$, while $C$ stores the set of candidate keys for the virtual points. We use $U$ to hold the loss for each candidate virtual point and vector $V$ to store the final optimal virtual points.
First, the algorithm identifies the sub-sequences of candidate virtual points that could have their minimal loss at the endpoints or in-between the sequence. This is shown in Lines 4 to 12. If there are more than two points in a sub-sequence, the candidate virtual point with the minimal loss can be within that sub-sequence. As such, the two endpoints of the sub-sequence are saved in array $G$ for calculating the partial derivatives.
Afterwards, in Lines 13 to 22, the partial derivative of the loss function with respect to the candidate virtual points is calculated for all point pairs in $G$ using the equations derived above. If the signs of the partial derivatives corresponding to the two endpoints of a sub-sequence are different (i.e., on opposite sides of the $x$-axis), the two endpoints are added to array $M$ for calculating the minimum point. As shown in Fig. 4, for the sub-sequences that contain candidate virtual points with minimal loss, the partial derivatives of the end points will appear on the two sides of the $x$-axis. These minimal points are added to array $C$ after they are calculated. If the partial derivatives of the two endpoints are on the same side of the $x$-axis, the minima is at one of the endpoints, as such the two end points are added to $C$.
Finally, Lines 23 to 31 compute the loss for each candidate virtual point in $C$ and select the point with the minimum loss, as long as the new loss is smaller than the existing loss obtained so far over $K$ and any previously inserted virtual points. This process is repeated until at most $\lambda$ virtual points are inserted, or when the loss is not reduced any further. When the algorithm terminates, the final virtual points in $V$ are returned.
Complexity analysis. Our proposed CDF smoothing algorithm reduces the computation of the loss over $K$ to just once, which takes $O(n)$ time. This process is repeated to find $\lambda$ optimal candidate virtual points. However, there is no need to recalculate the loss function after adding a virtual point, as we could treat the key set with the previous virtual point inserted as the new original or base key set for a constant time calculation. Thereby, giving a time complexity of $O(\lambda+n)$.
5. CDF Smoothing for Hierarchical Indexes